### v1

# Fermion-to-Fermion Low-Density Parity-Check Codes

Abstract
Simulating fermionic systems on qubit-based quantum computers often demands significant computational resources due to the requirement to map fermions to qubits. Thus, designing a fault-tolerant quantum computer that operates directly with fermions offers an effective solution to this challenge. Here, we introduce a protocol for fault-tolerant fermionic quantum computation utilizing fermion-to-fermion low-density parity-check codes. Our method employs a fermionic low-density parity-check code memory, which transfers its state to fermionic color code processors assisted by lattice surgery, where logical operations are subsequently performed. To construct the fermionic low-density parity-check memory, we develop a systematic approach for creating fermionic stabilizer codes based on self-dual Calderbank-Shor-Steane codes. We present examples demonstrating these codes’ significantly improved coding rate compared to the fermionic color code. Finally, we simulate the dynamics of a fermionic system using our protocol, illustrating effective error suppression.
The study of strongly correlated fermionic systems is central to high-energy physics [1], material science [2], and quantum chemistry [3], promising insights into phenomena ranging from quark dynamics [4] to high-temperature superconductivity [5]. However, solving these problems using classical computers is often very challenging. For instance, quantum Monte Carlo methods usually face the sign problem when addressing fermionic problems [6]. In this context, quantum computers offer a promising alternative [7]. However, since conventional quantum computers use qubits, solving fermionic problems requires mapping fermionic operators to qubit operators, which often incurs substantial overhead [8, 9, 10, 11, 12, 13], making experimental implementation significantly challenging in the near term. To address this challenge, developing programmable fermionic quantum processors is increasingly appealing. In light of the unavoidable presence of noise, fault-tolerant fermionic quantum computing based on fermion-to-fermion repetition and color codes have been proposed very recently [14, 15]. However, the protocol’s overhead remains substantial because the encoding rate is low, with each block encoding only a single logical fermion.
Recently, protocols based on quantum low-density parity-check (qLDPC) codes have been proposed to reduce the overhead in qubit-based fault-tolerant quantum computation [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Compared to the paradigmatic surface code, qLDPC codes feature a significantly higher coding rate. Although these codes require long-range connectivity between qubits, recent technological advances in platforms [40, 41, 37, 42] such as Rydberg atom arrays, superconducting qubits, and trapped ions have made the near-term implementation of these codes promising. Despite these significant advancements, existing studies primarily focus on constructing fault-tolerant quantum computers with qubits. It remains unclear how to construct a fermion-to-fermion LDPC code capable of encoding multiple fermion modes in a single code block and how to perform logical operations on such codes.
Figure 1: Schematic illustration of fault-tolerant fermionic computation framework. The architecture comprises three components: fermionic memory implemented using fermionic LDPC codes, fermionic processor implemented using fermionic color codes to execute logical operations, and fermionic interface composed of ancilla fermions serving as a critical bridge for communication between memory and processor.
Here, we introduce a protocol for fault-tolerant fermionic quantum computation based on fermion-to-fermion LDPC codes. Our approach employs a fermionic LDPC memory integrated with fermionic color code processors as illustrated in Fig. 1, which is generalized from the qLDPC case. Logical fermionic information is initially stored in the fermionic LDPC memory and, when needed, is transferred to fermionic color code processors, where fault-tolerant logical operations are conducted. Upon completion, the information is returned to the memory. To construct the fermionic LDPC memory, we develop a systematic workflow for creating fermionic stabilizer codes based on self-dual Calderbank-Shor-Steane (CSS) codes. These codes are characterized by logical Majorana operators consisting of an odd number of physical Majorana operators. We construct three distinct classes of fermionic LDPC codes using three different LDPC codes: bicycle codes [43], finite Euclidean geometry codes [44, 45], and finite projective geometry codes [46]. These codes demonstrate a significantly improved coding rate compared to the fermionic color code. Furthermore, we propose a fault-tolerant method to transfer logical fermionic states between the fermionic memory and processor via lattice surgery [47, 48]. We show that this process maintains the code distance. Finally, we simulate the dynamics of a fermionic system, demonstrating effective error suppression.
We start by introducing the construction of our fermionic LDPC codes from $2n$ physical Majorana fermion operators $\left\{{\gamma}_{1},{\gamma}_{1}^{\prime},{\gamma}_{2},{\gamma}_{2}^{\prime},\dots,{\gamma}_{n},{\gamma}_{n}^{\prime}\right\}$ arising from $n$ physical complex fermion’s creation and annihilation operators [49]. The strings of these Majorana operators, together with a phase factor $\eta\in\left\{\pm 1,\pm\text{i}\right\}$, generate the group of Majorana operators $\text{Maj}(2n)\equiv\left\{{\Gamma}=\eta\prod_{j=1}^{n}{\gamma}_{j}^{\alpha_{j}}({\gamma}_{j}^{\prime})^{\alpha_{j}^{\prime}}|\alpha_{j},\alpha_{j}^{\prime}\in\{0,1\}\right\}$ [50]. The number of Majorana operators in $\Gamma$} is referred to as its weight. The Majorana stabilizer code is defined by a Majorana stabilizer group $\mathcal{S}_{\text{maj}}$, a subgroup of $\text{Maj}(2n)$, where all elements are Hermitian, mutually commute, ensuring it is an Abelian subgroup, have even weight to preserve the parity of a physical fermion system, and I-I with $I$ being the identity element in the group is excluded [50]. The logical operators are generated by an independent subset of $\text{Maj}(2n)$ comprising Majorana operator strings that commute with all elements of the stabilizer group $\mathcal{S}_{\text{maj}}$ but are not members of $\mathcal{S}_{\text{maj}}$ [51].
Each Majorana string operator can be represented by a binary vector $(\alpha_{1},\alpha_{1}^{\prime},\dots,\alpha_{n},\alpha_{n}^{\prime})$, where $\alpha_{j},\alpha_{j}^{\prime}\in\{0,1\}$ and $1\leq j\leq n$. The binary vectors corresponding to the $m$ independent stabilizer generators of $\mathcal{S}_{\text{maj}}$ form an $m\times n$ check matrix. For a fermion-to-fermion LDPC code, we construct a check matrix $H$ based on a self-dual CSS code with the binary check matrices that satisfy $H_{X}=H_{Z}=A$ and $AA^{T}=0$, as follows,
$H=\left(\begin{matrix}H_{\gamma}&0\\ 0&H_{\gamma^{\prime}}\end{matrix}\right),$ (1)
where $H_{\gamma}=H_{\gamma^{{}^{\prime}}}=A$. $H_{\gamma}$ and $H_{\gamma^{\prime}}$ correspond to check matrices that characterize the stabilizers consisting exclusively of $\{\gamma_{j}\}$ and $\{\gamma_{j}^{\prime}\}$ operators, respectively, as described for fermion-to-fermion color codes in Ref. [14]. The logical operators are represented by vectors in the kernal of $H$ that cannot be generated by stabilizers, forming a homology group $\ker(A)/\text{im}(A)$ for both $\gamma$ and $\gamma^{\prime}$ types of logical operators [52].
In our protocol, there are multiple logical fermions within a single fermionic LDPC memory, and these logical fermions are transferred to processors for the execution of logical operations. It is essential that the logical Majorana operator in memory anticommutes with that in the processor to adhere to fermionic statistics. We thus require all logical Majorana operators in memory have odd weight and even overlap with each other [50]. Such conditions also ensure that the fermionic statistics are maintained when concatenating different fermionic code blocks [14]. We denote the logical Majorana operators as $\overline{\gamma}_{j}$ and $\overline{\gamma}_{j}^{\prime}$ with $1\leq j\leq k$ ($k$ is the number of logical complex fermions), which consist of physical Majorana operators $\{\gamma_{1},\dots,\gamma_{n}\}$ and $\{\gamma_{1}^{\prime},\dots,\gamma_{n}^{\prime}\}$, respectively. The logical complex fermion creation operators are then defined as $\overline{c}_{j}^{\dagger}=\frac{1}{2}(\overline{\gamma}_{j}+\text{i}\overline{\gamma}_{j}^{\prime})$ [49].
Figure 2: (a) Error model used to benchmark the capacity of our fermionic LDPC code memory. Each sample consists of 1010 cycles, each incorporating three layers: random errors, syndrome measurements, and error correction. (b)-(d) The numerically calculated logical failure rate with respect to the physical error rate for three fermionic LDPC codes: $\text{PG}(2,3)$, $\text{EG}(2,3)$, and Bicycle code $100,20,10]]_{\mathrm{f}}$. The dashed red line indicates a power law fit to the data, and the gray line represents the case where the logical and physical error rates are equal. The vertical dashed gray line highlights the code’s pseudo-threshold.
Figure 3: (a) Quantum circuit to transfer a logical fermionic state $|\overline{\psi}\rangle_{A}$ to code block $B$ (initialized as $|\overline{0}\rangle_{B}$). The protocol involves: (i) measuring $i\overline{\gamma}_{A}\overline{\gamma}_{B}$ and applying the $Z$ gate of $\exp(i\pi\overline{n}_{A})$, provided the measurement result is 1-1; (ii) measuring the logical particle number $\overline{n}_{A}$ and applying the $D$ gate defined as $\overline{D}=\exp(i\frac{\pi}{2}\overline{\gamma})$ to both code block $A$ and $B$ if the measurement outcome is 11. The validation of this circuit is provided in Supplemental material. (b) Illustration of the lattice surgery for fault-tolerant measurements of $i\overline{\gamma}_{A}\overline{\gamma}_{B}$ for a case with $|\overline{\gamma}_{A}|=|\overline{\gamma}_{B}|$. In this example, code $B$ is a $d=5$ fermionic color code, and code $A$ is a fermionic LDPC code containing five Majorana operators in the support of $\overline{\gamma}_{A}$. The empty circles represent the $\gamma$-type Majorana operators in blocks AA and $B$. Several pairs of ancilla Majorana operators are introduced with each horizontal pair represented by yellow and gray circles to denote $\gamma$-type and $\gamma^{\prime}$-type Majorana operator forming a complex fermion. The $\gamma$-type stabilizer generators at the boundary, connected to ancilla fermions, are modified to incorporate the ancilla Majorana operators (green rhombuses). Measurement and gauge stabilizer generators are depicted as blue squares and red triangles, respectively. The product of these stabilizer generators yields the joint logical operator $i\overline{\gamma}_{A}\overline{\gamma}_{B}$. The $\gamma^{\prime}$-type Majorana stabilizers remain unchanged. The general procedure for introducing ancilla Majorana operators and constructing the corresponding stabilizers is detailed in Supplemental Material. (d) A more general fermionic lattice surgery scheme for two fermionic LDPC codes with arbitrary weights.
Previous studies have primarily focused on encoding Majorana modes into qubits, lacking methods for identifying odd-weight logical operators within codes [50, 53, 54]. To identify them, we generalize the Gram-Schmidt orthogonalization process to vector spaces over $\mathbb{F}_{2}$ [55]. Let $\left\{\vec{\gamma}_{j}|j=1,\dots,k^{\prime},k^{\prime}\geq k\right\}$ be a basis for $\ker{A}/\text{im}(A)$, which is itself also a vector space; $\vec{\gamma}_{j}$ represents the corresponding logical $\overline{\gamma}_{j}$ or $\overline{\gamma}^{\prime}_{j}$ operator, expressed as $\gamma_{1}^{[\vec{\gamma}_{j}]_{1}}\dots\gamma_{n}^{[\vec{\gamma}_{j}]_{n}}$ or $(\gamma_{1}^{\prime})^{[\vec{\gamma}_{j}]_{1}}\dots(\gamma_{n}^{\prime})^{[\vec{\gamma}_{j}]_{n}}$. If there exists a basis vector with odd number of $1^{\prime}$s, say $\vec{\gamma}_{1}$, we remove it from the basis and make the remaining basis vectors orthogonal to it via $\vec{\gamma}_{j}\to\vec{\gamma}_{j}-(\vec{\gamma}_{1}\cdot\vec{\gamma}_{j})\vec{\gamma}_{1}$, i.e., $\vec{\gamma}_{1}$ has even overlap with all others. This procedure is repeated until no odd-weight basis vector remains. The obtained odd-weight vectors correspond to logical operators which have odd weight, and the overlap weight between any two of them is even, thus satisfying the fermionic anticommutation relation. The necessary and sufficient condition for the existence of at least one odd-weight logical is $(1,1,\dots,1)^{T}\notin\text{im}(A)$ (see Supplemental Material). We also prove that the procedure always produces a maximumally linearly independent set of odd-weight binary vectors (see Supplemental Material). In this case, the resulting odd-weight logical operators define a fermionic subspace code for our fermionic LDPC code. In addition, we provide a necessary and sufficient condition under which all basis vectors can be used to generate odd-weight logical operators in Supplemental Material.
Based on three different LDPC codes, which include finite projective geometry codes [46], finite Euclidean geometry codes [44, 45], and bicycle codes [43], we construct three classes of fermionic LDPC codes (see Supplemental Material for detailed construction methods). We find that these subspace fermionic LDPC codes yield the same or nearly the same encoding rate as the original ones. We consider one code from each type and benchmark their logical information resilience as a fermion memory through numerical simulations. Specifically, we initialize the logical state in $|\overline{0}\overline{0}\dots\overline{0}\rangle$, a logical fermionic vacuum state without any logical fermions, and simulate errors occurring as physical single-fermion gate ${\gamma}_{j}$, ${\gamma}_{j}^{\prime}$, and $i{\gamma}_{j}{\gamma}_{j}^{\prime}$ with equal probability $p/3$, where $p$ denotes the physical error rate. Subsequently, we perform error detection via syndrome measurements (assuming perfect syndrome measurements for simplicity), followed by error decoding using the belief propagation and ordered-statistical decoding (BP+OSD) algorithm [56, 57]. After decoding, we correct errors by applying physical single-fermion gates and then repeat the entire process for the corrected new state. For each sample, we conduct $N_{c}=10$ rounds of error-correction circuit cycles (each cycle contains three layers: random errors, syndrome measurements, and error correction), as shown in Fig. 2(a). The logical error probability for each sample is defined as $P_{\mathrm{L}}(N_{c})=N_{\mathrm{error}}/N_{\mathrm{sample}}$, where $N_{\mathrm{error}}$ and $N_{\mathrm{sample}}$ denote the number of erroneous trials and total trials, respectively. The logical failure rate is then $p_{\mathrm{L}}=1-(1-P_{\mathrm{L}}(N_{c}))^{1/N_{c}}$ [36, 37]. Figure 2 shows the logical error rate with respect to the physical error rate for the finite projective geometry $\text{PG}(3)$ code, the finite Euclidean geometry $\text{EG}(2,4)$ code, and the $100,20,10]]_{\mathrm{f}}$ bicycle code. We clearly see that the logical error rate is significantly suppressed compared to the physical one according to $p_{\mathrm{L}}\sim p^{\alpha}$ with $\alpha\approx 2$. Following the convention [37], we define the code’s pseudo-threshold as the solution to the break-even equation $p_{\mathrm{L}}(p)=P(p,k)$, where $P(p,k)$ is the probability of at least one error occurring on $k$ physical fermion sites at physical error rate $p$. Under this error model, all three codes exhibit relatively high pseudo-thresholds.
In Supplemental Material, we also study several other self-dual codes such as Kitaev Majorana code [49] and unicycle codes [43]. We find that they are not suitable for producing well-behaved fermionic LDPC because they do not support the generation of odd-weight logical operators.
We now study how to perform logical operations on the fermionic LDPC memory. Instead of executing logical gates directly on the fermionic LDPC, we consider a method whereby the states of a target logical fermion mode in memory are transferred to an external fermionic color code processor. Logical operations are then performed on these processors, and the states are subsequently transferred back to the memory. This approach is inspired by fault-tolerant measurements of logical operators in qubit LDPC codes [36, 30, 58]. If the logical operation involves two logical fermion modes, we consider two color code processors. The logical information of the two fermion modes is individually transferred, logical operations are applied between these two color code blocks, and the information is subsequently returned to memory. Logical operations involving more logical fermion modes can be performed similarly. For fermionic state transfers between memory and processor, we design a measurement-based circuit as illustrated in Fig. 3(a). This circuit enables the transfer of a single logical fermion information described by logical Majorana operators $\overline{\gamma}_{A}$ and $\overline{\gamma}_{A}^{\prime}$ to a processor described by logical Majorana operators $\overline{\gamma}_{B}$ and $\overline{\gamma}_{B}^{\prime}$, which is initialized in the logical vacuum state $|\overline{0}\rangle_{B}$. In other words, an output state is the same as the initial state except that the processor logical fermion plays the role of the corresponding logical fermion in memory. During the transfer, a key step is the fault-tolerant measurement of the joint logical operator $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$, as shown in Fig. 3(a).
Figure 4: (a) The quantum circuit used to benchmark our scheme. We first initialize the system in the logical state $\left|\overline{0}\overline{1}\overline{0}\overline{1}\right\rangle$ and then apply three braid gates with each gate executing the operation ${B}_{j,j+1}=\exp(i\pi(\overline{c}_{j}^{\dagger}\overline{c}_{j+1}+\text{H.c.})/2)$. Onsite measurements of $\overline{n}_{j}=\overline{c}_{j}^{\dagger}\overline{c}_{j}$ with $j=0,1,2$ are then performed, followed by a fermionic swap (fSWAP) gate if the measurement result is zero. The unitary gates and feedback process constitute a module, which is executed consecutively multiple times. (b) The expectation value of $\overline{n}_{0}$ with respect to time which is characterized by the number of executed modules. We see that the simulation under noise with error correction (filled blue circles) is very close to the noiseless simulation result (black dashed line), in stark contrast to the simulation results under noise without error correction (filled green circles).
We propose fermionic lattice surgery to realize the fault-tolerant joint measurement as shown in Fig. 3(b) and (c). Let $d_{A}$ be the weight of $\overline{\gamma}_{A}$ and $\overline{\gamma}_{A}^{\prime}$. For simplicity, we consider the case where block $B$ is a color code whose distance satisfies that $d_{B}=d_{A}$ and put the general case where block $B$ is also a fermionic LDPC code in Supplemental Material. To realize the measurement, we align the support of the logical operator $\overline{\gamma}_{A}$ and $\overline{\gamma}_{B}$ and introduce a set $Q_{C}$ of $4(d_{A}-1)$ ancilla Majorana operators (corresponding to $2(d_{A}-1)$ complex fermions) between these two codes, labeled as $\gamma_{a,1},\gamma_{a,1}^{\prime},\dots,\gamma_{a,(d_{A}-1)},\gamma_{a,(d_{A}-1)}^{\prime}$ and $\gamma_{b,1},\gamma_{b,1}^{\prime},\dots,\gamma_{b,(d_{A}-1)},\gamma_{b,(d_{A}-1)}^{\prime}$. The original $\gamma$-type stabilizer generators at the boundary are modified by including two ancilla Majorana operators as shown in Fig. 3(b). Measurement stabilizer generators $M_{1},M_{2},\dots,M_{d_{A}}$ and gauge stabilizer generators $G_{1},G_{2},\dots,G_{d_{A}-1}$ are also introduced. The original $\gamma^{\prime}$-type stabilizer generators for both block of $A$ and $B$ are not modified. This merging procedure creates a merged stabilizer code $\mathcal{C}_{\mathrm{merged}}$ where $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$ becomes a stabilizer. Since $2(d_{A}-1)$ ancilla complex fermions and $2d_{A}-1$ new stabilizers are introduced, this new code only contains one logical complex fermion described by $\overline{\gamma}_{A}^{\prime}$ and $\overline{\gamma}_{B}^{\prime}$. In other words, one cannot find $\gamma$-type logical operators in the merged code. In this case, we show that the merging procedure does not decrease the code distance based on the formalism of the subsystem code (see Supplemental Material). The number of the Majorana modes introduced during the whole process is small compared to the fermionic LDPC memory, resulting in a small resource overhead (see Supplemental Material). The Supplemental Material also provides the general guidelines for introducing ancilla Majorana operators and constructing stabilizer generators. Moreover, we propose a more comprehensive scheme for performing lattice surgery between two fermionic stabilizer codes (see Fig. 3(c) for an example) with details in Supplemental Material.
To achieve the fault-tolerant measurement of $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$, we first initialize the physical ancilla complex fermions in the $|0\rangle$ state and then measure all the stabilizer generators in the merged code and perform $\min\{d_{A},d_{B}\}$ round of error corrections in the presence of noisy measurements to ensure fault-tolerance. Finally, each physical ancilla fermion is measured in the particle number basis to return the state to the original code space.
We now demonstrate the capability of our codes by simulating the quantum circuit of fermions shown in Fig. 4(a). The circuit incorporates braid operations that enable the tunneling of fermions between neighboring sites, and onsite measurements, followed by fermionic swap gates, provided the measurement outcome is zero. For an initial state $|\overline{0}\overline{1}\overline{0}\overline{1}\rangle$, we expect that the steady state of this dynamical process is characterized by the particle occupation number at the upper boundary exceeding the average particle occupation number of $1/2$ due to feedback effects [59, 60, 61]. To simulate the dynamic behavior, we utilize the $\text{PG}(3)$ code as our fermionic LDPC code and select four encoded complex fermion modes from it. Additionally, we introduce two copies of the fermionic Steane code to function as processors. The calculated time evolution of the particle number at the first site, denoted $\left\langle\hat{n}_{0}\right\rangle$, is shown in Fig. 4(b). The value of $\left\langle\hat{n}_{0}\right\rangle$ exhibits oscillatory behavior, but the time-averaged value is larger than $1/2$, which aligns with the expected characteristics of a skin state. Notably, with our fault-tolerant logical operations and error correction techniques, the error-corrected data demonstrate significantly higher fidelity compared to the uncorrected data, as shown in Fig. 4(b). The time evolution results using the general lattice surgery is provided in Supplemental Material.
In summary, we have developed a systematic workflow for constructing fermionic LDPC code memory based on self-dual CSS codes and proposed methods for executing logical operations based on lattice surgery. We demonstrate that our fermionic LDPC code can be used to simulate dynamics of fermionic systems fault-tolerantly with effective error detection and correction. Such fermion-to-fermion quantum computation has the potential to significantly reduce the computational complexity associated with simulating fermionic systems, leveraging the possibility of achieving quantum advantage in quantum simulations. Our work encourages the pursuit of high-performance fermionic LDPC codes with higher code rates and larger code distances, as well as the advancement of more efficient methods for executing logical operations with reduced overhead.
Note added. During the final stage of this work, we became aware of a related work [62], which uses Majorana LDPC codes to encode logical qubits. In contrast, we encode logical fermions with Majorana LDPC codes, and our fermionic state transfer protocol provides an alternative method to perform logical operations with low overhead.
Acknowledgements.
We thank Y. Li, Q. Xu, and F. Wei for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12474265 and No. 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604). We also acknowledge the support by center of high performance computing, Tsinghua University.

### v2

@@
 Thus, designing a fault-tolerant quantum computer that operates directly with fermions offers an effective solution to this challenge.
-Here, we introduce a protocol for fault-tolerant fermionic quantum computation utilizing fermion-to-fermion low-density parity-check codes.
-Our method employs a fermionic low-density parity-check code memory, which transfers its state to fermionic color code processors assisted by lattice surgery, where logical operations are subsequently performed.
-To construct the fermionic low-density parity-check memory, we develop a systematic approach for creating fermionic stabilizer codes based on self-dual Calderbank-Shor-Steane codes.
-We present examples demonstrating these codes’ significantly improved coding rate compared to the fermionic color code.
+Here, we introduce a protocol for fault-tolerant fermionic quantum computation utilizing fermion-to-fermion low-density parity-check (LDPC) codes.
+Our method employs a fermionic LDPC memory, which transfers its state to fermionic color code processors, where logical operations are subsequently performed.
+We propose using odd-weight logical Majorana operators to form the code space, serving as memory for the fermionic LDPC code, and provide an algorithm to identify these logical operators.
+We present examples showing that the coding rate of fermionic codes often matches that of qubit codes, while the logical error rate can be significantly lower than the physical error rate.
+Furthermore, we propose two methods for performing fermionic lattice surgery to facilitate state transfer.
 Finally, we simulate the dynamics of a fermionic system using our protocol, illustrating effective error suppression.
@@
 However, the protocol’s overhead remains substantial because the encoding rate is low, with each block encoding only a single logical fermion.
-Recently, protocols based on quantum low-density parity-check (qLDPC) codes have been proposed to reduce the overhead in qubit-based fault-tolerant quantum computation [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
-Compared to the paradigmatic surface code, qLDPC codes feature a significantly higher coding rate.
+Recently, protocols based on quantum LDPC codes have been proposed to reduce the overhead in qubit-based fault-tolerant quantum computation [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
+Compared to the paradigmatic surface code, quantum LDPC codes feature a significantly higher coding rate.
 Although these codes require long-range connectivity between qubits, recent technological advances in platforms [40, 41, 37, 42] such as Rydberg atom arrays, superconducting qubits, and trapped ions have made the near-term implementation of these codes promising.
@@
 Our approach employs a fermionic LDPC memory integrated with fermionic color code processors as illustrated in Fig.
-1, which is generalized from the qLDPC case.
+1, which is generalized from the quantum LDPC case [36].
 Logical fermionic information is initially stored in the fermionic LDPC memory and, when needed, is transferred to fermionic color code processors, where fault-tolerant logical operations are conducted.
@@
 These codes demonstrate a significantly improved coding rate compared to the fermionic color code.
-Furthermore, we propose a fault-tolerant method to transfer logical fermionic states between the fermionic memory and processor via lattice surgery [47, 48].
+Furthermore, we propose two methods for performing fermionic lattice surgery to facilitate transfer of logical fermionic states between the fermionic memory and processor.
 We show that this process maintains the code distance.
@@
 The logical complex fermion creation operators are then defined as $\overline{c}_{j}^{\dagger}=\frac{1}{2}(\overline{\gamma}_{j}+\text{i}\overline{\gamma}_{j}^{\prime})$ [47].
-Figure 2: (a) Error model used to benchmark the capacity of our fermionic LDPC code memory.
+Figure 2: (a) Error model used to benchmark the capacity of our fermionic LDPC code memory initialized in either the logical vacuum state $\overline{|{0}\dots 0\rangle}$ or the logical product state $\overline{|{+}\dots+\rangle}$, where $\overline{\gamma}\overline{|{+}\rangle}=\overline{|{+}\rangle}$.
 Each sample consists of 1010 cycles, each incorporating three layers: random errors, syndrome measurements, and error correction.
-(b)-(d) The numerically calculated logical failure rate with respect to the physical error rate for three fermionic LDPC codes: $\text{PG}(2,3)$, $\text{EG}(2,3)$, and Bicycle code $100,20,10]]_{\mathrm{f}}$.
-The dashed red line indicates a power law fit to the data, and the gray line represents the case where the logical and physical error rates are equal.
+(b)-(d) The numerically calculated logical failure rate with respect to the physical error rate for three fermionic LDPC codes: $\text{PG}(2,8)$, $\text{EG}(2,8)$, and Bicycle code $100,20,7]]_{\mathrm{f}}$.
+The dashed red line indicates a power law fit to the data, and the gray line represents the probability that an error occurs on at least one physical site (see Supplemental Material).
 The vertical dashed gray line highlights the code’s pseudo-threshold.
 Figure 3: (a) Quantum circuit to transfer a logical fermionic state $|\overline{\psi}\rangle_{A}$ to code block $B$ (initialized as $|\overline{0}\rangle_{B}$).
-The protocol involves: (i) measuring $i\overline{\gamma}_{A}\overline{\gamma}_{B}$ and applying the $Z$ gate of $\exp(i\pi\overline{n}_{A})$, provided the measurement result is 1-1; (ii) measuring the logical particle number $\overline{n}_{A}$ and applying the $D$ gate defined as $\overline{D}=\exp(i\frac{\pi}{2}\overline{\gamma})$ to both code block $A$ and $B$ if the measurement outcome is 11.
+The protocol involves: (i) measuring $i\overline{\gamma}_{A}\overline{\gamma}_{B}$ and applying the $Z$ gate of $\exp(i\pi\overline{n}_{B})$, provided the measurement result is 1-1; (ii) measuring the logical particle number $\overline{n}_{A}$ and applying the $D$ gate defined as $\overline{D}=\exp(i\frac{\pi}{2}\overline{\gamma})$ to both code block $A$ and $B$ if the measurement outcome is 11; (iii) applying a phase rotation gate to the state in the processor.
 The validation of this circuit is provided in Supplemental material.
-(b) Illustration of the lattice surgery for fault-tolerant measurements of $i\overline{\gamma}_{A}\overline{\gamma}_{B}$ for a case with $|\overline{\gamma}_{A}|=|\overline{\gamma}_{B}|$.
+(b) Illustration of the first method for lattice surgery achieving fault-tolerant measurements of $i\overline{\gamma}_{A}\overline{\gamma}_{B}$ in the case with $|\overline{\gamma}_{A}|=|\overline{\gamma}_{B}|$.
 In this example, code $B$ is a $d=5$ fermionic color code, and code $A$ is a fermionic LDPC code containing five Majorana operators in the support of $\overline{\gamma}_{A}$.
-The empty circles represent the $\gamma$-type Majorana operators in blocks AA and $B$.
+The empty circles represent the $\gamma$-type Majorana operators in blocks $A$ and $B$.
 Several pairs of ancilla Majorana operators are introduced with each horizontal pair represented by yellow and gray circles to denote $\gamma$-type and $\gamma^{\prime}$-type Majorana operator forming a complex fermion.
@@
 The general procedure for introducing ancilla Majorana operators and constructing the corresponding stabilizers is detailed in Supplemental Material.
-(d) A more general fermionic lattice surgery scheme for two fermionic LDPC codes with arbitrary weights.
+(d) The second method for fermionic lattice surgery between two fermionic LDPC codes with arbitrary weights.
 Previous studies have primarily focused on encoding Majorana modes into qubits, lacking methods for identifying odd-weight logical operators within codes [48, 51, 52].
@@
 The obtained odd-weight vectors correspond to logical operators which have odd weight, and the overlap weight between any two of them is even, thus satisfying the fermionic anticommutation relation.
-The necessary and sufficient condition for the existence of at least one odd-weight logical is $(1,1,\dots,1)^{T}\notin\text{im}(A)$ (see Supplemental Material).
-We also prove that the procedure always produces a maximumally linearly independent set of odd-weight binary vectors (see Supplemental Material).
-In this case, the resulting odd-weight logical operators define a fermionic subspace code for our fermionic LDPC code.
-In addition, we provide a necessary and sufficient condition under which all basis vectors can be used to generate odd-weight logical operators in Supplemental Material.
-Based on three different LDPC codes, which include finite projective geometry codes [46], finite Euclidean geometry codes [44, 45], and bicycle codes [43], we construct three classes of fermionic LDPC codes (see Supplemental Material for detailed construction methods).
-We find that these subspace fermionic LDPC codes yield the same or nearly the same encoding rate as the original ones.
+The necessary and sufficient condition for the existence of at least one odd-weight logical is $(1,1,\dots,1)^{T}\notin\text{im}(A)$ [48].
+We find that for most cases we consider the procedure produces a linearly independent set containing $k^{\prime}$ odd-weight binary vectors (see Supplemental Material).
+The resulting odd-weight logical operators define a fermionic subspace code for our fermionic LDPC code.
+Based on three different LDPC codes, which include finite projective geometry codes [46], finite Euclidean geometry codes [44, 45], and bicycle codes [43], we construct three classes of fermionic LDPC codes (see Supplemental Material).
+We find that these subspace fermionic LDPC codes yield the same encoding rate as the original ones.
 We consider one code from each type and benchmark their logical information resilience as a fermion memory through numerical simulations.
-Specifically, we initialize the logical state in $|\overline{0}\overline{0}\dots\overline{0}\rangle$, a logical fermionic vacuum state without any logical fermions, and simulate errors occurring as physical single-fermion gate ${\gamma}_{j}$, ${\gamma}_{j}^{\prime}$, and $i{\gamma}_{j}{\gamma}_{j}^{\prime}$ with equal probability $p/3$, where $p$ denotes the physical error rate.
+Specifically, we randomly initialize the logical state in either the logical vacuum state $\overline{|{0}\dots 0\rangle}$ or the logical product state $\overline{|{+}\dots+\rangle}$, where $\overline{\gamma}\overline{|{+}\rangle}=\overline{|{+}\rangle}$, and simulate errors occurring as physical single-fermion gate ${\gamma}_{j}$, ${\gamma}_{j}^{\prime}$, and $i{\gamma}_{j}{\gamma}_{j}^{\prime}$ with equal probability $p/3$, where $p$ denotes the physical error rate.
 Subsequently, we perform error detection via syndrome measurements (assuming perfect syndrome measurements for simplicity), followed by error decoding using the belief propagation and ordered-statistical decoding (BP+OSD) algorithm [54, 55].
@@
 The logical failure rate is then $p_{\mathrm{L}}=1-(1-P_{\mathrm{L}}(N_{c}))^{1/N_{c}}$ [36, 37].
-Figure 2 shows the logical error rate with respect to the physical error rate for the finite projective geometry $\text{PG}(3)$ code, the finite Euclidean geometry $\text{EG}(2,4)$ code, and the $100,20,10]]_{\mathrm{f}}$ bicycle code.
+Figure 2 shows the logical error rate with respect to the physical error rate for the finite projective geometry $\text{PG}(2,8)$ code, the finite Euclidean geometry $\text{EG}(2,4)$ code, and the $100,20,7]]_{\mathrm{f}}$ bicycle code.
 We clearly see that the logical error rate is significantly suppressed compared to the physical one according to $p_{\mathrm{L}}\sim p^{\alpha}$ with $\alpha\approx 2$.
@@
 Logical operations are then performed on these processors, and the states are subsequently transferred back to the memory.
-This approach is inspired by fault-tolerant measurements of logical operators in qubit LDPC codes [36, 30, 58].
+This approach is inspired by fault-tolerant measurements of logical operators in qubit LDPC codes [56, 57, 36, 30, 58].
 If the logical operation involves two logical fermion modes, we consider two color code processors.
@@
 The unitary gates and feedback process constitute a module, which is executed consecutively multiple times.
-(b) The expectation value of $\overline{n}_{0}$ with respect to time which is characterized by the number of executed modules.
-We see that the simulation under noise with error correction (filled blue circles) is very close to the noiseless simulation result (black dashed line), in stark contrast to the simulation results under noise without error correction (filled green circles).
-We propose fermionic lattice surgery to realize the fault-tolerant joint measurement as shown in Fig.
+(b) The expectation value of logical particle numbers at site 11 and $4$ with respect to time which is characterized by the number of executed modules.
+We see that the simulation under noise with error correction (filled blue circles and squares) is very close to the noiseless simulation result (blue and red dotted-dashed lines), in stark contrast to the simulation results under noise without error correction (filled green circles and squares).
+The error bars for blue circles and green squares are hidden behind the symbols.
+We propose two methods for fermionic lattice surgery to realize the fault-tolerant joint measurement as shown in Fig.
 3(b) and (c).
 Let $d_{A}$ be the weight of $\overline{\gamma}_{A}$ and $\overline{\gamma}_{A}^{\prime}$.
-For simplicity, we consider the case where block $B$ is a color code whose distance satisfies that $d_{B}=d_{A}$ and put the general case where block $B$ is also a fermionic LDPC code in Supplemental Material.
+For simplicity, we consider the case where code $B$ is a fermionic color code whose distance satisfies that $d_{B}=d_{A}$ and put the general case where code $B$ is also a fermionic LDPC code in Supplemental Material.
 To realize the measurement, we align the support of the logical operator $\overline{\gamma}_{A}$ and $\overline{\gamma}_{B}$ and introduce a set $Q_{C}$ of $4(d_{A}-1)$ ancilla Majorana operators (corresponding to $2(d_{A}-1)$ complex fermions) between these two codes, labeled as $\gamma_{a,1},\gamma_{a,1}^{\prime},\dots,\gamma_{a,(d_{A}-1)},\gamma_{a,(d_{A}-1)}^{\prime}$ and $\gamma_{b,1},\gamma_{b,1}^{\prime},\dots,\gamma_{b,(d_{A}-1)},\gamma_{b,(d_{A}-1)}^{\prime}$.
@@
 In other words, one cannot find $\gamma$-type logical operators in the merged code.
-In this case, we show that the merging procedure does not decrease the code distance based on the formalism of the subsystem code (see Supplemental Material).
+In this case, we show that the merging procedure does not decrease the code distance based on the formalism of the subsystem code [59, 60] (see Supplemental Material).
 The number of the Majorana modes introduced during the whole process is small compared to the fermionic LDPC memory, resulting in a small resource overhead (see Supplemental Material).
-The Supplemental Material also provides the general guidelines for introducing ancilla Majorana operators and constructing stabilizer generators.
+The Supplemental Material also provides the guidelines for introducing ancilla Majorana operators and constructing stabilizer generators for this method.
 Moreover, we propose a more comprehensive scheme for performing lattice surgery between two fermionic stabilizer codes (see Fig.
@@
 The circuit incorporates braid operations that enable the tunneling of fermions between neighboring sites, and onsite measurements, followed by fermionic swap gates, provided the measurement outcome is zero.
-For an initial state $|\overline{0}\overline{1}\overline{0}\overline{1}\rangle$, we expect that the steady state of this dynamical process is characterized by the particle occupation number at the upper boundary exceeding the average particle occupation number of $1/2$ due to feedback effects [59, 60, 61].
-To simulate the dynamic behavior, we utilize the $\text{PG}(3)$ code as our fermionic LDPC code and select four encoded complex fermion modes from it.
+For an initial state $|\overline{0}\overline{1}\overline{0}\overline{1}\rangle$, we expect that the steady state of this dynamical process is a skin state where fermionic particles mainly reside in the upper half part due to feedback effects [61, 62, 63].
+To simulate the dynamic behavior, we utilize the $\text{PG}(2,8)$ code as our fermionic LDPC code and select four encoded complex fermion modes from it.
 Additionally, we introduce two copies of the fermionic Steane code to function as processors.
-The calculated time evolution of the particle number at the first site, denoted $\left\langle\hat{n}_{0}\right\rangle$, is shown in Fig.
+The calculated time evolution of the particle number at the first and fourth site is shown in Fig.
 4(b).
-The value of $\left\langle\hat{n}_{0}\right\rangle$ exhibits oscillatory behavior, but the time-averaged value is larger than $1/2$, which aligns with the expected characteristics of a skin state.
+The value of $\left\langle\overline{n}_{1}\right\rangle$ suddenly increases to one while the value of $\left\langle\overline{n}_{4}\right\rangle$ decreases to zero, which aligns with the expected characteristics of a skin state.
 Notably, with our fault-tolerant logical operations and error correction techniques, the error-corrected data demonstrate significantly higher fidelity compared to the uncorrected data, as shown in Fig.
@@
 We demonstrate that our fermionic LDPC code can be used to simulate dynamics of fermionic systems fault-tolerantly with effective error detection and correction.
+Given that our protocol employs gates like the $D$ gate, which does not conserve the parity of fermions, one can consider using referenced fermionic modes as our physical modes, which can be realized with the assistance of auxiliary reference modes [15].
 Such fermion-to-fermion quantum computation has the potential to significantly reduce the computational complexity associated with simulating fermionic systems, leveraging the possibility of achieving quantum advantage in quantum simulations.

### v3

@@
 We propose using odd-weight logical Majorana operators to form the code space, serving as memory for the fermionic LDPC code, and provide an algorithm to identify these logical operators.
-We present examples showing that the coding rate of fermionic codes often matches that of qubit codes, while the logical error rate can be significantly lower than the physical error rate.
+We present examples showing that the encoding rate of fermionic codes often matches that of qubit codes, while the logical error rate can be significantly lower than the physical error rate.
 Furthermore, we propose two methods for performing fermionic lattice surgery to facilitate state transfer.
@@
 However, since conventional quantum computers use qubits, solving fermionic problems requires mapping fermionic operators to qubit operators, which often incurs substantial overhead [8, 9, 10, 11, 12, 13], making experimental implementation significantly challenging in the near term.
-To address this challenge, developing programmable fermionic quantum processors is increasingly appealing.
+To address this challenge, developing programmable fermionic quantum processors is increasingly appealing [10, 14].
 In light of the unavoidable presence of noise, fault-tolerant fermionic quantum computing based on fermion-to-fermion repetition and color codes have been proposed very recently [15, 16].
@@
 Recently, protocols based on quantum LDPC codes have been proposed to reduce the overhead in qubit-based fault-tolerant quantum computation [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
-Compared to the paradigmatic surface code, quantum LDPC codes feature a significantly higher coding rate.
+Compared to the paradigmatic surface code, quantum LDPC codes feature a significantly higher encoding rate.
 Although these codes require long-range connectivity between qubits, recent technological advances in platforms [41, 42, 38, 43] such as Rydberg atom arrays, superconducting qubits, and trapped ions have made the near-term implementation of these codes promising.
@@
 It remains unclear how to construct a fermion-to-fermion LDPC code capable of encoding multiple fermion modes in a single code block and how to perform logical operations on such codes.
-Figure 1: Schematic illustration of fault-tolerant fermionic computation framework.
+Figure 1: Schematic illustration of fault-tolerant fermionic quantum computation framework.
 The architecture comprises three components: fermionic memory implemented using fermionic LDPC codes, fermionic processor implemented using fermionic color codes to execute logical operations, and fermionic interface composed of ancilla fermions serving as a critical bridge for communication between memory and processor.
@@
 We construct three distinct classes of fermionic LDPC codes using three different LDPC codes: bicycle codes [44], finite Euclidean geometry codes [45, 46], and finite projective geometry codes [47].
-These codes demonstrate a significantly improved coding rate compared to the fermionic color code.
+These codes can have a significantly higher encoding rate compared to the fermionic color code.
 Furthermore, we propose two methods for performing fermionic lattice surgery to facilitate transfer of logical fermionic states between the fermionic memory and processor.
@@
 The general procedure for introducing ancilla Majorana operators and constructing the corresponding stabilizers is detailed in Supplemental Material.
-(d) The second method for fermionic lattice surgery between two fermionic LDPC codes with arbitrary weights.
-Previous studies have primarily focused on encoding Majorana modes into qubits, lacking methods for identifying odd-weight logical operators within codes [48, 51, 52].
+(d) Illustration of the second method for fermionic lattice surgery between two fermionic LDPC codes with arbitrary weights.
+Previous studies have primarily focused on encoding Majorana modes into qubits [49, 52, 53].
+While some results also demonstrate the existence of odd-weight logical operators in Majorana surface codes [54], there is a lack of methods for systematically identifying these operators within a fermonic LDPC code.
 To identify them, we generalize the Gram-Schmidt orthogonalization process to vector spaces over $\mathbb{F}_{2}$ [55].
@@
 We consider one code from each type and benchmark their logical information resilience as a fermion memory through numerical simulations.
-Specifically, we randomly initialize the logical state in either the logical vacuum state $\overline{|{0}\dots 0\rangle}$ or the logical product state $\overline{|{+}\dots+\rangle}$, where $\overline{\gamma}\overline{|{+}\rangle}=\overline{|{+}\rangle}$, and simulate errors occurring as physical single-fermion gate ${\gamma}_{j}$, ${\gamma}_{j}^{\prime}$, and $i{\gamma}_{j}{\gamma}_{j}^{\prime}$ with equal probability $p/3$, where $p$ denotes the physical error rate.
+Specifically, we randomly initialize the logical state in either the logical vacuum state $\overline{|{0}\dots 0\rangle}$ or the logical product state $\overline{|{+}\dots+\rangle}$ ($\overline{\gamma}\overline{|{+}\rangle}=\overline{|{+}\rangle}$), both of which can be prepared by a series of braiding operations [56, 57].
+We then simulate errors occurring as physical single-fermion gate ${\gamma}_{j}$, ${\gamma}_{j}^{\prime}$, and $i{\gamma}_{j}{\gamma}_{j}^{\prime}$ with equal probability $p/3$, where $p$ denotes the physical error rate.
 Subsequently, we perform error detection via syndrome measurements (assuming perfect syndrome measurements for simplicity), followed by error decoding using the belief propagation and ordered-statistical decoding (BP+OSD) algorithm [58, 59].
@@
 3(a).
-Figure 4: (a) The quantum circuit used to benchmark our scheme.
+Figure 4: (a) Quantum circuit used to benchmark our protocol.
 We first initialize the system in the logical state $\left|\overline{0}\overline{1}\overline{0}\overline{1}\right\rangle$ and then apply three braid gates with each gate executing the operation ${B}_{j,j+1}=\exp(i\pi(\overline{c}_{j}^{\dagger}\overline{c}_{j+1}+\text{H.c.})/2)$.
@@
 4(b).
-The time evolution results using the general lattice surgery is provided in Supplemental Material.
+The time evolution results using the second method for lattice surgery is provided in Supplemental Material.
 In summary, we have developed a systematic workflow for constructing fermionic LDPC code memory based on self-dual CSS codes and proposed methods for executing logical operations based on lattice surgery.
@@
 Given that our protocol employs gates like the $D$ gate, which does not conserve the parity of fermions, one can consider using referenced fermionic modes as our physical modes, which can be realized with the assistance of auxiliary reference modes [16].
-Such fermion-to-fermion quantum computation has the potential to significantly reduce the computational complexity associated with simulating fermionic systems, leveraging the possibility of achieving quantum advantage in quantum simulations.
-Our work encourages the pursuit of high-performance fermionic LDPC codes with higher code rates and larger code distances, as well as the advancement of more efficient methods for executing logical operations with reduced overhead.
-Note added.
-During the final stage of this work, we became aware of a related work [64], which uses Majorana LDPC codes to encode logical qubits.
-In contrast, we encode logical fermions with Majorana LDPC codes, and our fermionic state transfer protocol provides an alternative method to perform logical operations with low overhead.
+Such fermion-to-fermion quantum computation has the potential to significantly reduce the computational complexity associated with simulating fermionic systems.
+Our work encourages the pursuit of high-performance fermionic LDPC codes with higher encoding rates and larger code distances, as well as the advancement of more efficient methods for executing logical operations with reduced overhead.
+Note added. During the final stage of this work, we became aware of a related work [68], which uses Majorana LDPC codes to encode logical qubits.
 Acknowledgements.

### v4

@@
 We propose using odd-weight logical Majorana operators to form the code space, serving as memory for the fermionic LDPC code, and provide an algorithm to identify these logical operators.
-We present examples showing that the encoding rate of fermionic codes often matches that of qubit codes, while the logical error rate can be significantly lower than the physical error rate.
+We present examples showing that the encoding rate of fermionic codes often matches that of qubit codes, while the logical failure rate can be significantly lower than the physical error rate.
 Furthermore, we propose two methods for performing fermionic lattice surgery to facilitate state transfer.
@@
 Compared to the paradigmatic surface code, quantum LDPC codes feature a significantly higher encoding rate.
-Although these codes require long-range connectivity between qubits, recent technological advances in platforms [41, 42, 38, 43] such as Rydberg atom arrays, superconducting qubits, and trapped ions have made the near-term implementation of these codes promising.
+Although these codes require long-range connectivity between qubits, recent technological advances in platforms [13, 12, 17, 22, 74] such as Rydberg atom arrays, superconducting qubits, and trapped ions have made the near-term implementation of these codes promising.
 Despite these significant advancements, existing studies primarily focus on constructing fault-tolerant quantum computers with qubits.
@@
 Upon completion, the information is returned to the memory.
-To construct the fermionic LDPC memory, we develop a systematic workflow for creating fermionic stabilizer codes based on self-dual Calderbank-Shor-Steane (CSS) codes.
+To construct the fermionic LDPC memory, we develop a systematic workflow for creating fermionic stabilizer codes based on weakly-self-dual Calderbank-Shor-Steane (CSS) codes.
 These codes are characterized by logical Majorana operators consisting of an odd number of physical Majorana operators.
-We construct three distinct classes of fermionic LDPC codes using three different LDPC codes: bicycle codes [44], finite Euclidean geometry codes [45, 46], and finite projective geometry codes [47].
+We construct four distinct classes of fermionic LDPC codes using four different weakly-self-dual LDPC codes: bicycle codes [46], finite Euclidean geometry codes [21, 5], finite projective geometry codes [25] and stacked LDPC codes [45].
 These codes can have a significantly higher encoding rate compared to the fermionic color code.
@@
 Finally, we simulate the dynamics of a fermionic system, demonstrating effective error suppression.
-We start by introducing the construction of our fermionic LDPC codes from $2n$ physical Majorana fermion operators $\left\{{\gamma}_{1},{\gamma}_{1}^{\prime},{\gamma}_{2},{\gamma}_{2}^{\prime},\dots,{\gamma}_{n},{\gamma}_{n}^{\prime}\right\}$ arising from $n$ physical complex fermion’s creation and annihilation operators [48].
+We start by introducing the construction of our fermionic LDPC codes from $2n$ physical Majorana fermion operators $\left\{{\gamma}_{1},{\gamma}_{1}^{\prime},{\gamma}_{2},{\gamma}_{2}^{\prime},\dots,{\gamma}_{n},{\gamma}_{n}^{\prime}\right\}$ arising from $n$ physical complex fermions’ creation and annihilation operators [36].
 The strings of these Majorana operators, together with a phase factor $\eta\in\left\{\pm 1,\pm\text{i}\right\}$, generate the group of Majorana operators $\text{Maj}(2n)\equiv\left\{{\Gamma}=\eta\prod_{j=1}^{n}{\gamma}_{j}^{\alpha_{j}}({\gamma}_{j}^{\prime})^{\alpha_{j}^{\prime}}|\alpha_{j},\alpha_{j}^{\prime}\in\{0,1\}\right\}$ [16].
@@
 Each Majorana string operator can be represented by a binary vector $(\alpha_{1},\alpha_{1}^{\prime},\dots,\alpha_{n},\alpha_{n}^{\prime})$, where $\alpha_{j},\alpha_{j}^{\prime}\in\{0,1\}$ and $1\leq j\leq n$.
-The binary vectors corresponding to the $m$ independent stabilizer generators of $\mathcal{S}_{\text{maj}}$ form an $m\times n$ check matrix.
-For a fermion-to-fermion LDPC code, we construct a check matrix $H$ based on a self-dual CSS code with the binary check matrices that satisfy $H_{X}=H_{Z}=A$ and $AA^{T}=0$, as follows,
+The binary vectors corresponding to the $m$ independent stabilizer generators of $\mathcal{S}_{\text{maj}}$ form an $m\times 2n$ check matrix.
+For a fermion-to-fermion LDPC code, we construct a check matrix $H$ based on a weakly-self-dual CSS code with the binary check matrices that satisfy $H_{X}=H_{Z}=A$ and $AA^{T}=0$, as follows,
 $H=\left(\begin{matrix}H_{\gamma}&0\\ 0&H_{\gamma^{\prime}}\end{matrix}\right),$ (1)
-where $H_{\gamma}=H_{\gamma^{{}^{\prime}}}=A$.
+where $H_{\gamma}=H_{\gamma^{\prime}}=A$.
 $H_{\gamma}$ and $H_{\gamma^{\prime}}$ correspond to check matrices that characterize the stabilizers consisting exclusively of $\{\gamma_{j}\}$ and $\{\gamma_{j}^{\prime}\}$ operators, respectively, as described for fermion-to-fermion color codes in Ref.
 [62].
-The logical operators are represented by vectors in the kernal of $H$ that cannot be generated by stabilizers, forming a homology group $\ker(A)/\text{im}(A)$ for both $\gamma$ and $\gamma^{\prime}$ types of logical operators [51].
-In our protocol, there are multiple logical fermions within a single fermionic LDPC memory, and these logical fermions are transferred to processors for the execution of logical operations.
+The logical operators are represented by vectors in the kernel of $H$ that cannot be generated by stabilizers, whose span forms a homology group $\ker(A)/\text{im}(A^{T})$ for both $\gamma$ and $\gamma^{\prime}$ types of logical operators [14].
+In our protocol, there are multiple logical fermion modes within a single fermionic LDPC memory, and these logical fermion modes are transferred to processors for the execution of logical operations.
 It is essential that the logical Majorana operator in memory anticommutes with that in the processor to adhere to fermionic statistics.
 We thus require all logical Majorana operators in memory have odd weight and even overlap with each other [16].
-Such conditions also ensure that the fermionic statistics are maintained when concatenating different fermionic code blocks [15].
-We denote the logical Majorana operators as $\overline{\gamma}_{j}$ and $\overline{\gamma}_{j}^{\prime}$ with $1\leq j\leq k$ ($k$ is the number of logical complex fermions), which consist of physical Majorana operators $\{\gamma_{1},\dots,\gamma_{n}\}$ and $\{\gamma_{1}^{\prime},\dots,\gamma_{n}^{\prime}\}$, respectively.
+Such conditions also ensure that the fermionic statistics are maintained when concatenating different fermion code blocks [62].
+We denote the logical Majorana operators as $\overline{\gamma}_{j}$ and $\overline{\gamma}_{j}^{\prime}$ with $1\leq j\leq k$ ($k$ is the number of logical complex fermion modes), which consist of physical Majorana operators $\{\gamma_{1},\dots,\gamma_{n}\}$ and $\{\gamma_{1}^{\prime},\dots,\gamma_{n}^{\prime}\}$, respectively.
 The logical complex fermion creation operators are then defined as $\overline{c}_{j}^{\dagger}=\frac{1}{2}(\overline{\gamma}_{j}+\text{i}\overline{\gamma}_{j}^{\prime})$ [36].
-Figure 2: (a) Error model used to benchmark the capacity of our fermionic LDPC code memory initialized in either the logical vacuum state $\overline{|{0}\dots 0\rangle}$ or the logical product state $\overline{|{+}\dots+\rangle}$, where $\overline{\gamma}\overline{|{+}\rangle}=\overline{|{+}\rangle}$.
-Each sample consists of 1010 cycles, each incorporating three layers: random errors, syndrome measurements, and error correction.
-(b)-(d) The numerically calculated logical failure rate with respect to the physical error rate for three fermionic LDPC codes: $\text{PG}(2,8)$, $\text{EG}(2,8)$, and Bicycle code $100,20,7]]_{\mathrm{f}}$.
-The dashed red line indicates a power law fit to the data, and the gray line represents the probability that an error occurs on at least one physical site (see Supplemental Material).
-The vertical dashed gray line highlights the code’s pseudo-threshold.
+Figure 2: (a) Illustration of the circuit-level noise model where the syndrome extraction transfer circuits used to establish entanglement between data fermions and ancilla qubits are noisy.
+The initial states are physical $|0\rangle$ states for fermions.
+In the end, a transversal measurement on all physical fermion modes in the particle number basis is performed.
+Logical failure rate $p_{L}$ versus physical error rate $p$ under (b) phenomenological and (c) circuit-level noise model for fermionic $20,4,4]]_{\text{f}}$, $36,4,6]]_{\text{f}}$ and $44,4,7]]_{\text{f}}$ double-chain bicycle codes.
+We also perform a power law fit to the data according to the fitting formula $p_{L}\sim p^{\alpha}$ with $\alpha=2.21,3.06,4.3$ in (b) and $\alpha=2.12,3.46,4.54$ in (c), and the gray lines represent the probability that at least an error occurs on four physical fermion modes.
+Here, the Tesseract decoder [2] is employed, and the results obtained using the belief propagation and ordered-statistical decoding (BP+OSD) [41, 59] algorithm are presented in Supplemental Material.
 Figure 3: (a) Quantum circuit to transfer a logical fermionic state $|\overline{\psi}\rangle_{A}$ to code block $B$ (initialized as $|\overline{0}\rangle_{B}$).
-The protocol involves: (i) measuring $i\overline{\gamma}_{A}\overline{\gamma}_{B}$ and applying the $Z$ gate of $\exp(i\pi\overline{n}_{B})$, provided the measurement result is 1-1; (ii) measuring the logical particle number $\overline{n}_{A}$ and applying the $D$ gate defined as $\overline{D}=\exp(i\frac{\pi}{2}\overline{\gamma})$ to both code block $A$ and $B$ if the measurement outcome is 11; (iii) applying a phase rotation gate to the state in the processor.
+The protocol involves: (i) measuring $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$ and applying the $Z$ gate of $\exp(\text{i}\pi\overline{n}_{B})$, provided the measurement result is 1-1; (ii) measuring the logical particle number $\overline{n}_{A}$ and applying two braiding gates and one fermionic $Z$ gate, if the measurement outcome is 11; (iii) applying a phase rotation $S$ gate to the state in the processor.
 The validation of this circuit is provided in Supplemental Material.
-(b) Illustration of the first method for lattice surgery achieving fault-tolerant measurements of $i\overline{\gamma}_{A}\overline{\gamma}_{B}$ in the case with $|\overline{\gamma}_{A}|=|\overline{\gamma}_{B}|$.
+(b) Illustration of the first method for lattice surgery achieving fault-tolerant measurements of $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$ in the case with $|\overline{\gamma}_{A}|=|\overline{\gamma}_{B}|$.
 In this example, code $B$ is a $d=5$ fermionic color code, and code $A$ is a fermionic LDPC code containing five Majorana operators in the support of $\overline{\gamma}_{A}$.
@@
 Measurement and gauge stabilizer generators are depicted as blue squares and red triangles, respectively.
-The product of these stabilizer generators yields the joint logical operator $i\overline{\gamma}_{A}\overline{\gamma}_{B}$.
+The product of these stabilizer generators yields the joint logical operator $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$.
 The $\gamma^{\prime}$-type Majorana stabilizers remain unchanged.
-The general procedure for introducing ancilla Majorana operators and constructing the corresponding stabilizers is detailed in Supplemental Material.
-(d) Illustration of the second method for fermionic lattice surgery between two fermionic LDPC codes with arbitrary weights.
 Previous studies have primarily focused on encoding Majorana modes into qubits [16, 43, 72].
-While some results also demonstrate the existence of odd-weight logical operators in Majorana surface codes [54], there is a lack of methods for systematically identifying these operators within a fermonic LDPC code.
+While some results also demonstrate the existence of odd-weight logical operators in Majorana surface codes [48], there is a lack of methods for systematically identifying these operators within a fermionic LDPC code.
 To identify them, we generalize the Gram-Schmidt orthogonalization process to vector spaces over $\mathbb{F}_{2}$ [57].
-Let $\left\{\vec{\gamma}_{j}|j=1,\dots,k^{\prime},k^{\prime}\geq k\right\}$ be a basis for $\ker{A}/\text{im}(A)$, which is itself also a vector space; $\vec{\gamma}_{j}$ represents the corresponding logical $\overline{\gamma}_{j}$ or $\overline{\gamma}^{\prime}_{j}$ operator, expressed as $\gamma_{1}^{[\vec{\gamma}_{j}]_{1}}\dots\gamma_{n}^{[\vec{\gamma}_{j}]_{n}}$ or $(\gamma_{1}^{\prime})^{[\vec{\gamma}_{j}]_{1}}\dots(\gamma_{n}^{\prime})^{[\vec{\gamma}_{j}]_{n}}$.
+Let $\left\{\vec{\gamma}_{j}|j=1,\dots,k^{\prime},k^{\prime}\geq k\right\}$ be a basis for $\ker{A}/\text{im}(A^{T})$, which is itself also a vector space; $\vec{\gamma}_{j}$ represents the corresponding logical $\overline{\gamma}_{j}$ or $\overline{\gamma}^{\prime}_{j}$ operator, expressed as $\gamma_{1}^{[\vec{\gamma}_{j}]_{1}}\dots\gamma_{n}^{[\vec{\gamma}_{j}]_{n}}$ or $(\gamma_{1}^{\prime})^{[\vec{\gamma}_{j}]_{1}}\dots(\gamma_{n}^{\prime})^{[\vec{\gamma}_{j}]_{n}}$.
 If there exists a basis vector with odd number of $1^{\prime}$s, say $\vec{\gamma}_{1}$, we remove it from the basis and make the remaining basis vectors orthogonal to it via $\vec{\gamma}_{j}\to\vec{\gamma}_{j}-(\vec{\gamma}_{1}\cdot\vec{\gamma}_{j})\vec{\gamma}_{1}$, i.e., $\vec{\gamma}_{1}$ has even overlap with all others.
@@
 The obtained odd-weight vectors correspond to logical operators which have odd weight, and the overlap weight between any two of them is even, thus satisfying the fermionic anticommutation relation.
-The necessary and sufficient condition for the existence of at least one odd-weight logical is $(1,1,\dots,1)^{T}\notin\text{im}(A)$ [49].
-We find that for most cases we consider the procedure produces a linearly independent set containing $k^{\prime}$ odd-weight binary vectors (see Supplemental Material).
-The resulting odd-weight logical operators define a fermionic subspace code for our fermionic LDPC code.
-Based on three different LDPC codes, which include finite projective geometry codes [47], finite Euclidean geometry codes [45, 46], and bicycle codes [44], we construct three classes of fermionic LDPC codes (see Supplemental Material).
-We find that these subspace fermionic LDPC codes yield the same encoding rate as the original ones.
-We consider one code from each type and benchmark their logical information resilience as a fermion memory through numerical simulations.
-Specifically, we randomly initialize the logical state in either the logical vacuum state $\overline{|{0}\dots 0\rangle}$ or the logical product state $\overline{|{+}\dots+\rangle}$ ($\overline{\gamma}\overline{|{+}\rangle}=\overline{|{+}\rangle}$), both of which can be prepared by a series of braiding operations [56, 57].
-We then simulate errors occurring as physical single-fermion gate ${\gamma}_{j}$, ${\gamma}_{j}^{\prime}$, and $i{\gamma}_{j}{\gamma}_{j}^{\prime}$ with equal probability $p/3$, where $p$ denotes the physical error rate.
-Subsequently, we perform error detection via syndrome measurements (assuming perfect syndrome measurements for simplicity), followed by error decoding using the belief propagation and ordered-statistical decoding (BP+OSD) algorithm [58, 59].
-After decoding, we correct errors by applying physical single-fermion gates and then repeat the entire process for the corrected new state.
-For each sample, we conduct $N_{c}=10$ rounds of error-correction circuit cycles (each cycle contains three layers: random errors, syndrome measurements, and error correction), as shown in Fig.
+The necessary and sufficient condition for the existence of at least one odd-weight logical is $(1,1,\dots,1)^{T}\notin\text{im}(A^{T})$ [16].
+We find that for all cases we consider, the procedure produces a linearly independent set containing $k^{\prime}$ odd-weight binary vectors (see Supplemental Material).
+Based on four different LDPC codes, which include finite projective geometry codes [25], finite Euclidean geometry codes [21, 5], bicycle codes [46], and stacked LDPC codes [45], we construct four classes of fermionic LDPC codes, which yield the same code parameters $n,k,d]]_{\text{f}}$ as the original qubit codes (see Supplemental Material).
+We now focus on three fermionic double-chain bicycle codes, which belong to the fermionic stacked LDPC codes, and evaluate their logical information resilience as a fermion memory through numerical simulations.
+Specifically, we initialize all physical fermion modes in the vacuum state, $|0\rangle$.
+The performance of a code depends on quantum error models employed.
+For completeness, we consider three types of error models, analogous to the qubit case: code-capacity, phenomenological and circuit-level error models (see Fig.
+2(a)) [70].
+The code-capacity error model does not account for errors in syndrome measurements and thus represents the upper bound of code performance against the errors considered.
+In the phenomenological and circuit-level error models, the syndrome measurement circuits are noisy: In the former, only the measurement outcomes are noisy, whereas in the latter, all operations in the circuit are also noisy (see Supplemental Material).
+Under these two error models, we perform $N_{c}=d$ rounds of syndrome measurements, followed by a transversal measurement on all physical fermion modes in the particle number basis, as illustrated in Fig.
 2(a).
-The logical error probability for each sample is defined as $P_{\mathrm{L}}(N_{c})=N_{\mathrm{error}}/N_{\mathrm{sample}}$, where $N_{\mathrm{error}}$ and $N_{\mathrm{sample}}$ denote the number of erroneous trials and total trials, respectively.
-The logical failure rate is then $p_{\mathrm{L}}=1-(1-P_{\mathrm{L}}(N_{c}))^{1/N_{c}}$ [37, 38].
-Figure 2 shows the logical error rate with respect to the physical error rate for the finite projective geometry $\text{PG}(2,8)$ code, the finite Euclidean geometry $\text{EG}(2,4)$ code, and the $100,20,7]]_{\mathrm{f}}$ bicycle code.
-We clearly see that the logical error rate is significantly suppressed compared to the physical one according to $p_{\mathrm{L}}\sim p^{\alpha}$ with $\alpha\approx 2$.
-Following the convention [38], we define the code’s pseudo-threshold as the solution to the break-even equation $p_{\mathrm{L}}(p)=P(p,k)$, where $P(p,k)$ is the probability of at least one error occurring on $k$ physical fermion sites at physical error rate $p$.
-Under this error model, all three codes exhibit relatively high pseudo-thresholds.
-In Supplemental Material, we also study several other self-dual codes such as Kitaev Majorana code [48] and unicycle codes [44].
-We find that they are not suitable for producing well-behaved fermionic LDPC because they do not support the generation of odd-weight logical operators.
+We then apply a decoder to interpret the obtained syndromes and verify whether logical errors occur.
+In numerical experiments, we perform $N_{\mathrm{sample}}$ trials and identify $N_{\mathrm{error}}$ erroneous occurrences.
+Consequently, the logical error probability is given by $P_{\mathrm{L}}(N_{c})=N_{\mathrm{error}}/N_{\mathrm{sample}}$.
+Conventionally, the logical failure rate is defined as $p_{\mathrm{L}}=1-(1-P_{\mathrm{L}}(N_{c}))^{1/N_{c}}\approx P_{L}/N_{c}$ [77, 17], which approximately characterizes the logical error rate per syndrome cycle.
+Figure 2(c) and (d) show the logical failure rate with respect to the physical error rate $p$ for three fermionic double-chain bicycle codes.
+We clearly see that the logical failure rate is significantly suppressed compared to the physical one according to $p_{\mathrm{L}}\sim p^{\alpha}$ with $\alpha\approx\lceil\frac{d}{2}\rceil$.
+By convention [17], we define the code’s pseudo-threshold as the solution to the break-even equation $p_{\mathrm{L}}(p)=P(p,k)$, where $P(p,k)=1-(1-p)^{k}$ is the probability of at least one error occurring on $k$ physical fermion modes at physical error rate $p$.
+Under the circuit-level error model, the calculated pseudo-thresholds of the $36,4,6]]_{\text{f}}$ and $44,4,7]]_{\text{f}}$ codes exceed $0.1\%$.
+In Supplemental Material, we also provide the numerical simulation results for fermionic finite projective geometry codes, fermionic finite Euclidean geometry codes, fermionic bicycle codes, fermionic double-layer bivariate bicycle (BB) codes, fermionic double-layer twisted BB codes, and fermionic double-layer reflection codes.
+In addition, we study several other self-dual codes such as Kitaev Majorana code [36] and unicycle codes [46] and find that they are not suitable for producing well-behaved fermionic LDPC because they do not support the generation of odd-weight logical operators.
 We now study how to perform logical operations on the fermionic LDPC memory.
@@
 This approach is inspired by fault-tolerant measurements of logical operators in qubit LDPC codes [35, 28, 77, 23, 56].
-If the logical operation involves two logical fermion modes, we consider two color code processors.
-The logical information of the two fermion modes is individually transferred, logical operations are applied between these two color code blocks, and the information is subsequently returned to memory.
-Logical operations involving more logical fermion modes can be performed similarly.
+If the logical operation involves two logical fermionic modes, we consider two color code processors.
+The logical information of the two fermionic modes is individually transferred, logical operations are applied between these two color code blocks, and the information is subsequently returned to memory.
+Logical operations involving more logical fermionic modes can be performed similarly.
 For fermionic state transfers between memory and processor, we design a measurement-based circuit as illustrated in Fig.
@@
 Figure 4: (a) Quantum circuit used to benchmark our protocol.
-We first initialize the system in the logical state $\left|\overline{0}\overline{1}\overline{0}\overline{1}\right\rangle$ and then apply three braid gates with each gate executing the operation ${B}_{j,j+1}=\exp(i\pi(\overline{c}_{j}^{\dagger}\overline{c}_{j+1}+\text{H.c.})/2)$.
+We first initialize the system in the logical state $\left|\overline{0}\overline{1}\overline{0}\overline{1}\right\rangle$ and then apply three tunneling gates with each gate executing the operation ${T}_{j,j+1}=\exp(\text{i}\pi(\overline{c}_{j}^{\dagger}\overline{c}_{j+1}+\text{H.c.})/2)$.
 Onsite measurements of $\overline{n}_{j}=\overline{c}_{j}^{\dagger}\overline{c}_{j}$ with $j=0,1,2$ are then performed, followed by a fermionic swap (fSWAP) gate if the measurement result is zero.
@@
 (b) The expectation value of logical particle numbers at site 11 and $4$ with respect to time which is characterized by the number of executed modules.
-We see that the simulation under noise with error correction (filled blue circles and squares) is very close to the noiseless simulation result (blue and red dotted-dashed lines), in stark contrast to the simulation results under noise without error correction (filled green circles and squares).
-The error bars for blue circles and green squares are hidden behind the symbols.
-We propose two methods for fermionic lattice surgery to realize the fault-tolerant joint measurement as shown in Fig.
-3(b) and (c).
-Let $d_{A}$ be the weight of $\overline{\gamma}_{A}$ and $\overline{\gamma}_{A}^{\prime}$.
-For simplicity, we consider the case where code $B$ is a fermionic color code whose distance satisfies that $d_{B}=d_{A}$ and put the general case where code $B$ is also a fermionic LDPC code in Supplemental Material.
-To realize the measurement, we align the support of the logical operator $\overline{\gamma}_{A}$ and $\overline{\gamma}_{B}$ and introduce a set $Q_{C}$ of $4(d_{A}-1)$ ancilla Majorana operators (corresponding to $2(d_{A}-1)$ complex fermions) between these two codes, labeled as $\gamma_{a,1},\gamma_{a,1}^{\prime},\dots,\gamma_{a,(d_{A}-1)},\gamma_{a,(d_{A}-1)}^{\prime}$ and $\gamma_{b,1},\gamma_{b,1}^{\prime},\dots,\gamma_{b,(d_{A}-1)},\gamma_{b,(d_{A}-1)}^{\prime}$.
-The original $\gamma$-type stabilizer generators at the boundary are modified by including two ancilla Majorana operators as shown in Fig.
-3(b).
+We see that the simulation under circuit-level noise with error correction (filled blue circles and squares) is close to the noiseless simulation result (the dotted-dashed lines), in stark contrast to the simulation results under noise without error correction (filled red circles and squares).
+In the simulation, we apply the first method for lattice surgery.
+Results using the second method are provided in the Supplemental Material.
+Here, we set the physical error rate $p=10^{-4}$.
+The error bars are hidden behind the symbols.
+We propose two methods for fermionic lattice surgery to realize the fault-tolerant joint measurement, as shown in Fig.
+3(b), Fig.
+5, and Fig.
+6.
+The first method applies when code $B$ is a fermionic color code with a distance $d_{B}$ such that $d_{B}=d_{A}$, where $d_{A}=|\overline{\gamma}_{A}|=|\overline{\gamma}_{A}^{\prime}|$ with $|\dots|$ denoting the weight of a Majorana operator string.
+The second method is applicable when either $d_{B}=d_{A}$ or $d_{B}\neq d_{A}$, and when code $B$ is either a fermionic color code or a fermionic LDPC code.
+The procedure proceeds by introducing ancilla Majorana operators between the two codes, modifying the stabilizers associated with $\overline{\gamma}_{A}$ and $\overline{\gamma}_{B}$, and adding new stabilizers that merge $A$ and $B$ into a larger fermionic LDPC code $\mathcal{C}_{\text{merged}}$.
+Since the joint operator $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$ belongs to the stabilizer group $\mathcal{S}_{\text{merged}}$ of $\mathcal{C}_{\text{merged}}$, it can be measured fault-tolerantly by collecting the outcomes of the stabilizer measurements in $\mathcal{C}_{\text{merged}}$.
+Figure 3 provides an example of fermionic lattice surgery using method 1.
+In this example, we align the support of the logical operator $\overline{\gamma}_{A}$ and $\overline{\gamma}_{B}$ and introduce a set $Q_{C}$ of $4(d_{A}-1)$ ancilla Majorana operators (corresponding to $2(d_{A}-1)$ complex fermion modes) between these two codes, labeled as $\gamma_{a,1},\gamma_{a,1}^{\prime},\dots,\gamma_{a,(d_{A}-1)},\gamma_{a,(d_{A}-1)}^{\prime}$ and $\gamma_{b,1},\gamma_{b,1}^{\prime},\dots,\gamma_{b,(d_{A}-1)},\gamma_{b,(d_{A}-1)}^{\prime}$.
+The original $\gamma$-type stabilizer generators at the boundary are modified by including two ancilla Majorana operators.
 Measurement stabilizer generators $M_{1},M_{2},\dots,M_{d_{A}}$ and gauge stabilizer generators $G_{1},G_{2},\dots,G_{d_{A}-1}$ are also introduced.
@@
 This merging procedure creates a merged stabilizer code $\mathcal{C}_{\mathrm{merged}}$ where $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$ becomes a stabilizer.
-Since $2(d_{A}-1)$ ancilla complex fermions and $2d_{A}-1$ new stabilizers are introduced, this new code only contains one logical complex fermion described by $\overline{\gamma}_{A}^{\prime}$ and $\overline{\gamma}_{B}^{\prime}$.
-In other words, one cannot find $\gamma$-type logical operators in the merged code.
-In this case, we show that the merging procedure does not decrease the code distance based on the formalism of the subsystem code [63, 64] (see Supplemental Material).
-The number of the Majorana modes introduced during the whole process is small compared to the fermionic LDPC memory, resulting in a small resource overhead (see Supplemental Material).
-The Supplemental Material also provides the guidelines for introducing ancilla Majorana operators and constructing stabilizer generators for this method.
-Moreover, we propose a more comprehensive scheme for performing lattice surgery between two fermionic stabilizer codes (see Fig.
-3(c) for an example) with details in Supplemental Material.
-To achieve the fault-tolerant measurement of $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$, we first initialize the physical ancilla complex fermions in the $|0\rangle$ state and then measure all the stabilizer generators in the merged code and perform $\min\{d_{A},d_{B}\}$ round of error corrections in the presence of noisy measurements to ensure fault-tolerance.
+In End Matter, we provide guidelines for introducing ancilla Majorana operators and constructing stabilizer generators for both methods.
+In Supplemental Material, we show that the distance of dressed logical operators is not decreased based on the formalism of the subsystem code [73, 4].
+We also show that the number of the Majorana modes introduced during the whole process is small compared to the fermionic LDPC memory, resulting in a small resource overhead.
+To achieve the fault-tolerant measurement of $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$, we first initialize the physical ancilla complex fermions in the $|0\rangle$ state and then measure all the stabilizer generators in the merged code and perform $\min\{d_{A},d_{B}\}$ rounds of error corrections in the presence of noisy measurements to ensure fault-tolerance.
 Finally, each physical ancilla fermion is measured in the particle number basis to return the state to the original code space.
-We now demonstrate the capability of our codes by simulating the quantum circuit of fermions shown in Fig.
+We now demonstrate the capability of our codes by simulating the quantum circuit of four fermion modes shown in Fig.
 4(a).
-The circuit incorporates braid operations that enable the tunneling of fermions between neighboring sites, and onsite measurements, followed by fermionic swap gates, provided the measurement outcome is zero.
+The circuit incorporates tunneling operations that enable the tunneling of fermions between neighboring sites, and onsite measurements, followed by fermionic swap gates, provided the measurement outcome is zero.
 For an initial state $|\overline{0}\overline{1}\overline{0}\overline{1}\rangle$, we expect that the steady state of this dynamical process is a skin state where fermionic particles mainly reside in the upper half part due to feedback effects [75, 26, 44].
-To simulate the dynamic behavior, we utilize the $\text{PG}(2,8)$ code as our fermionic LDPC code and select four encoded complex fermion modes from it.
-Additionally, we introduce two copies of the fermionic Steane code to function as processors.
+To simulate the dynamical behavior, we employ the fermionic $88,4,7]]_{\text{f}}$ double-chain bicycle code as our fermionic LDPC code and use the four logical complex fermion modes without overlap for the simulation.
+Additionally, we introduce two copies of the fermionic $37,1,7]]_{\text{f}}$ color code to function as processors.
 The calculated time evolution of the particle number at the first and fourth site is shown in Fig.
@@
 4(b).
-The time evolution results using the second method for lattice surgery is provided in Supplemental Material.
-In summary, we have developed a systematic workflow for constructing fermionic LDPC code memory based on self-dual CSS codes and proposed methods for executing logical operations based on lattice surgery.
+In summary, we have developed a systematic workflow for constructing fermionic LDPC code memory based on weakly-self-dual CSS codes and proposed methods for executing logical operations based on lattice surgery.
 We demonstrate that our fermionic LDPC code can be used to simulate dynamics of fermionic systems fault-tolerantly with effective error detection and correction.
-Given that our protocol employs gates like the $D$ gate, which does not conserve the parity of fermions, one can consider using referenced fermionic modes as our physical modes, which can be realized with the assistance of auxiliary reference modes [16].
 Such fermion-to-fermion quantum computation has the potential to significantly reduce the computational complexity associated with simulating fermionic systems.
@@
 Li, Q.
-Xu, and F.
-Wei for helpful discussions.
+Xu, F.
+Wei, and Y.
+Mao for helpful discussions.
 This work is supported by the National Natural Science Foundation of China (Grant No.